A body starts from rest, accelerates uniformly. Find the ratio of displacement in, 
     (a) One, two and three seconds.
(b) First, second and third second.

Grade:11

1 Answers

Naitik jain
6 years ago
u=0 T =1Ut +1/2 at² =s1/2a1²=s11/2 a =s1T=21/2at²=S21/2 a2²=s21/2a4=S2T=31/2at²=s31/2a3²=s3S1:S2: S3= 1:3:9Distance travelled in first sec=1/2a (X) Distance travelled in 2nd secondS1-S2 1/2a4--1/2a1=1/2a3(y)Distance travelled in third second 1/2a9-1/2a4=1/2a5 (z )X:y:z=1:3:5